OpenStudy (anonymous):

domain of the function f(x)=3vx^2-x+1 it means cube root of x^2-x+1 domain of the function f(x)=3vx^2-x+1 it means cube root of x^2-x+1 @Mathematics

jimthompson5910 (jim_thompson5910):

Since you can take the cube root of any number, the domain is the set of all real numbers.

OpenStudy (anonymous):

how can i solve it out pls? my proff is very strict in marking

jimthompson5910 (jim_thompson5910):

there's not much to solve the cube root of any number is always defined also, the domain of any polynomial is the set of all real numbers, so this means that the radicand will always be defined. So this means that no matter what value of x you plug in, you'll always get some number as a result.

OpenStudy (anonymous):

which means -infinity to infinity?

jimthompson5910 (jim_thompson5910):

yep, the domain is \[\Large \left(-\infty,\infty\right)\]